An 80 kg object is dropped from rest from a height of 25 meters above the ground. Use the concept of energy conservation to find its speed when it is 10 meters above the ground.
Answer #1

Using law of conservation of energy,

0.5 mv^2 = mg (h1 - h2)

Velocity, v = sqrt(2 g (h1 - h2))

V = sqrt(2 x 9.8 x (25 - 10)) = 17.2 m/s

Comment in case any doubt please rate my answer...